a. The following information is from the national income accounts for country X



Y = C+I+G+(X-M)



C = 20+0.8Yd



T = 30



G = 22



X = 20



M = 4+0.3Y



Yd = Y-T



I=30



informs the above model, list all the endogenous and exogenous variables


Determine the equilibrium values for all the endogenous variables

1
Expert's answer
2021-11-07T19:39:59-0500

Y = C+I+G+(X-M)



C = 20+0.8Yd

This is the consumption function where 20 is autonomous consumption, 0.8 is the

marginal propensity to consume. and, Yd is disposable after tax income.

Yd=Y-T , where Y is national income and T is tax revenues.

Tax Revenues+0.3Y

0.3 is the average income tax rate

I=investment =30

G=Government spending =22

X=Exports =20

M=Imports=4+0.3Y

Step 1. Determine the aggregate expenditure function.

using the numbers above we get

Y = C+I+G+(X-M)

Y = 20+ 0.8Yd + 30 + 22 +(20-4+0.3Y)

Step 2. The equation for the 45-degree line is the set of points where GDP or national

income on the horizontal axis is equal to aggregate expenditure on the vertical axis

This equation becomes

Y= 20+ 0.8Yd + 30 + 22 +(20-4+0.3Y)

Y = 20 + 0.8 (Y-T) +30 + 22 + (-16 + 0.3y)

Step 3= insert the term T= 30 for the tax rate T. this produces an equation with only one variable Y.

variable, Y.

Y= 32+ 1.1 Y

Y-1.1Y= 32

-0.1Y=32

Y =-320